If three times a number added to 7 is divided by the number plus 9, the result is four thirds. Find the number.
Answer by jorel1380(3719)

Let x be your number:
7+3x/x+9=4/3
3(7+3x)=4(x+9)
21+9x=4x+36
9x-4x=36-21
5x=15
x=3